# Generative AI models enable efficient and physically consistent sea-ice simulations

### v1

@@ 2 Realistic long-term simulations through generative short-term predictions
GenSIM is influenced by long-term effects for the thermodynamics and dynamics as similarly found for the coupled neXtSIM-OPA simulation (Fig. 5) with strikingly well matching large-scale patterns. For the thermodynamics, GenSIM exhibits the same zones of ice growth and melting as neXtSIM (Fig. 5, a and b): main growing zones are found in the Laptev sea and Kara sea, as well as the Canadian archipelago and central Arctic, while the most pronounced melting zones are around Greenland, in the Hudson bay, the Barents sea, and the Chukchi sea. In all of these melting zones, sea ice is in contact with water coming from the Atlantic or Pacific ocean, connecting them to bottom melting (Fig. 5, c). The neural network correctly identifies these zones such that GenSIM exhibits there similar melting rates as neXtSIM-OPA. Hence, the learned long-term impact on the growth and melting in GenSIM is similar to the thermodynamical impact from the ocean.
@@
Although the dynamics are dominated by the atmosphere at short time scales [52], they expose a signature from the deeper ocean at longer time scales. Consequently, we find an alignment in the sea-ice drift and the velocity at the ocean surface in coupled sea-ice–ocean simulations (Fig. 5, e and f): sea ice circulates in front of the Alaskan and Canadian coast due to the Beaufort gyre, while the transpolar drift transports sea ice from growing zones to melting zones through the Fram strait. GenSIM’s sea-ice drift matches these major ocean current in the Arctic basin (Fig. 5, d), even though the effective Beaufort gyre is displaced. Moreover, when analysing the sea-ice speed near the Fram strait (Fig. 5, g), GenSIM’s drift has a similar evolution as neXtSIM’s. Notably, GenSIM shows the sea-ice fastening similar to neXtSIM from 2003 to 2008, which remains unexplained by atmospheric forcings but aligns to an increase in the ocean velocity.
@@
GenSIM has seemingly learned the long-term dynamics caused by ocean currents without having access to an external ocean forcing. Combined with the learned long-term thermodynamical patterns, we hypothesise that GenSIM has learned an internal ocean emulator based on the given sea-ice conditions and the external atmospheric forcings. Solely trained by the criterion to predict the tendencies over 12 hours, GenSIM correctly infers slow year-long effects and uncovers hidden dynamics, even when their signal-to-noise ratio is diminishing in the training data. These rather surprising results explain how GenSIM can match its long-term evolution to numerical models and satellite observations in a warming climate.
3 Consequences for AI-based Earth system models
@@
These results pave a way towards the use of generative diffusion and flow matching models for surrogates of other Earth system components than the atmosphere. For sea-ice modelling, they improve upon deterministic models in terms of forecasting scores but also in matching the observed energy spectra at all scales, while being able to produce large forecasting ensembles. These large ensembles are especially useful for sea-ice data assimilation, which has been hindered by high computational costs and the non-existence of adjoints [53]. Hence, generative surrogate models unlock a large potential to improve sea-ice forecasting in the future.
@@
In this study, we built upon the observation that generative flow models work similarly to classical numerical models: they integrate a dynamical system that is defined by the neural network. This similarity allows us to bridge methods used for classical numerical models to surrogate models. Here, we use domain decomposition to efficiently apply such models at global scale. Combined with the employed transformer architecture, the surrogate is flexible in its use with applications from regional to pan-Arctic scales. This flexibility enables us to train generative surrogates across heterogenous datasets and resolutions [32], potentially further boosting their performance.
@@
Our instantiated model, GenSIM, can extrapolate to long-term climate conditions, suggesting that generative surrogates can generalise short-term training objectives to long-term projections. We partially attribute this stability to GenSIM’s physical consistency: since the effects of the underlying physics are exposed by correlations that change over time, this physical consistency is a strength of generative diffusion and flow matching models, which excel in learning high-dimensional correlations across scales. Furthermore, when each autoregressive prediction time step is initialised from the same noise, GenSIM becomes unstable after the first iteration, indicating the importance of its stochastic generation process. Therefore, the long-term stability is likely an inherently emerging capability of generative flows employed for surrogate modelling.
@@
The successful extrapolation to long-term evolutions from short-term objectives emerges from learning the influence of slow components on the faster, targetted, time scales. This allows us for example to unveil the ocean impact on sea ice, even when GenSIM is solely forced by the atmosphere. Consequently, generative flow models can uncover hidden dynamics in the data, even for diminishing signal-to-noise ratios. Hence, by cycling their training with data assimilation [54, 55], generative surrogates have a huge potential to learn unknown physics from observations.
@@
The most common strategy for Earth system modelling is to have models for each Earth system component which interact through coupling. Yet, this strategy can fail for AI-based component models: as we have demonstrated, the dynamics of these models show an entanglement between components, here sea ice and the ocean. This makes it challenging to isolate the dynamics of a single component due to external forcings from other components. Therefore, we have to rethink how to do coupled Earth system modelling, when using AI-based models.
4 Methods
4.1 Data used for GenSIM
@@
The model is trained and compared to a 24-year-long simulation (1995-2018) [36] produced by neXtSIM [5], an advanced sea-ice model employing Lagrangian finite elements and a brittle Bingham-Maxwell rheology [6] (data available at https://github.com/sasip-climate/catalog-shared-data-SASIP/blob/main/outputs/OPA-neXtSIM_CREG025-ILBOXE140.md, last access: 18th August 2025). For this simulation, neXtSIM has been coupled to OPA, the ocean component of the NEMO modelling framework [37] and has been forced with the ERA5 reanalysis [45] for the atmosphere. The curvilinear mesh of our model corresponds to the Eulerian Arakawa-C mesh of OPA, to which the model output of neXtSIM has been interpolated. GenSIM predicts the sea-ice thickness, concentration, damage, $u$- and $v$-velocity, and the snow-on-ice thickness. All variables but the damage are averaged in a six hour window, while the damage is an instantaneous state.
@@
In addition to the initial conditions, GenSIM takes external atmospheric forcings from the ERA5 reanalysis dataset [45]. Originally on a latitude-longitude mesh with a $1/4^{\circ}$ resolution, the dataset is interpolated to our curvilinear mesh by nearest neighbour. The forcings consist of the 2-metre temperature, the 2-metre specific humidity, the 10-metre equatorial and meridional wind velocities. The specific humidity is estimated from the 2-metre temperature, 2-metre dew point, and surface pressure.
@@
To obtain neural network inputs that contain a memory about the past development without having access to the ocean, we construct degree day features from the 2-metre temperature [56, 57]. Using a temperature threshold of $T_{\textrm{thres}}=271.35\,\text{K}$, we construct the positive, $PDD$, and negative degree days, $NDD$, by
$PDD(T_{\mathrm{2m}},\Delta t)$ $=\int^{t}_{t-\Delta t}\mathrm{maximum(T_{\mathrm{2m}}(s)-T_{\textrm{thres}},0)}ds$
$\text{and }NDD(T_{\mathrm{2m}},\Delta t)$ $=\int^{t}_{t-\Delta t}\mathrm{minimum(T_{\mathrm{2m}}(s)-T_{\textrm{thres}},0)}ds,$
@@ 4.2 Evaluation data
We evaluate our long-term simulation with the sea-ice–ocean reanalysis product from the Pan-Arctic Ice-Ocean Modeling and Assimilation System (PIOMAS) [49] (data available at https://psc.apl.uw.edu/research/projects/arctic-sea-ice-volume-anomaly, last access: 18th August 2025). The here used dataset is on a one-degree latitudinal-longitudinal mesh and consists of monthly averaged sea-ice thickness and concentration values.
@@
The sea-ice volume is additionally evaluated with the CS2/SMOS product [50] (Version 2.6, data available at https://spaces.awi.de/spaces/CS2SMOS/overview, last access: 18th August 2025), which merges satellite observations from the CryoSat2 and Soil Moisture and Ocean Salinity (SMOS) missions. As CS2/SMOS infers solely the sea-ice thickness, it is complemented by sea-ice concentration observations of the EUMETSAT’s climate data record from the OSISAF project [58], which merges satellite measurements from SSMR, SSM/I, and SSMIS. Both datasets are on the EASE mesh with $25\,\text{km}$ resolution.
The deformation rates from the sea-ice dynamics are evaluated with the Lagrangian data from the RADARSAT Geophysical Processing System (RGPS) [59] (data available at https://asf.alaska.edu/datasets/daac/sea-ice-measures, last access: 18th August 2025). This dataset tracks features between two successive Synthetic Aperture Radar (SAR) images and provides high-resolution sea-ice motion data at a typical $10\,\text{km}$ spatial resolution and a time resolution of around three days.
4.3 Surrogate modelling with censored flow matching
@@
In the following, we expose the most important equations to understand how GenSIM and the deterministic model are trained. For a more rigorous treatment and a mathematical derivation of the equations, we refer to Appendix B.
@@
For a forecast window of 12 hours, based on the current state $\mathbf{x}_{t}$ and the forcings in the same window $\mathbf{F}_{t:t+12\,\mathrm{h}}$, the surrogate model should predict the sea-ice state $\mathbf{x}_{t+12\,\mathrm{h}}$–consisting of sea-ice thickness, concentration, damage, the velocities in $x$- and $y$-direction, and the snow-on-ice thickness. Instead of predicting the future state directly, the model outputs predicted tendencies $\widehat{\mathbf{z}}$, which are scaled by the per-variable climatological standard deviation of the tendencies $\boldsymbol{\sigma}_{\textrm{clim}}$ such that we can recover the prediction by
@@
$\widehat{\mathbf{x}}_{t+12\,\text{h}}=\mathbf{x}_{t}+\boldsymbol{\sigma}_{\textrm{clim}}\widehat{\mathbf{z}}$. (1)
@@
In our autoregressive framework, the subsequent prediction $\widehat{\mathbf{x}}_{t+24\,\text{h}}$ is performed with the previous prediction and updated forcings $\mathbf{F}_{t+12\,\mathrm{h}:t+24\,\mathrm{h}}$ as new inputs. This way, we can extend the prediction as long as possible [60, 28].
@@
For training, we have access to a training dataset with data triplets $(\mathbf{x}_{t},\mathbf{x}_{t+12\,\mathrm{h}},\mathbf{F}_{t:t+12\,\mathrm{h}})\sim\mathcal{D}$. From these data triplets, we can construct the targetted tendencies, $\mathbf{z}=(\mathbf{x}_{t+12\,\mathrm{h}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\textrm{clim}}$.
The deterministic model directly predicts the normalised tendencies with a neural network and its weights and biases $\boldsymbol{\theta}$, $\widehat{\mathbf{z}}_{\textrm{det}}=g_{\boldsymbol{\theta}}(\mathbf{x}_{t},\mathbf{F}_{t:t+12\,\mathrm{h}})$. The loss function of the deterministic model is then simply the mean-squared error between the output and the targetted tendencies,
$\mathcal{L}_{\textrm{det}}(\boldsymbol{\theta})=\mathbb{E}_{\mathcal{D}}\left[\lVert\mathbf{z}-g_{\boldsymbol{\theta}}(\mathbf{x}_{t},\mathbf{F}_{t:t+12\,\mathrm{h}})\rVert^{2}_{2}\right],$ (2)
where we take the norm over all variables and all unmasked grid points. As we work in normalised tendency space, the loss function automatically weights the different state variables by the reciprocal of their standard deviation, $1/(\boldsymbol{\sigma}_{\textrm{clim}})^{2}$, as similarly used for other deterministic surrogate models [60, 35].
@@
For GenSIM, we instantiate a generative flow to produce stochastic predictions [24, 48]. The generative flow is defined by the integration of an ordinary differential equation (ODE) in pseudo time $\tau\in[0,1]$,
$\frac{d\mathbf{z}_{\tau}}{d\tau}=\mathbf{u}$ (3)
@@
with its velocity $\mathbf{u}$ and the initial conditions at pseudo time $\tau=0$,
$\mathbf{z}_{0}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}).$ (4)
@@
Note that the subscript $\tau$ describes the dependency on the pseudo time, whereas $t$ is the real time.
@@
During training, when we know the targetted tendencies, we construct a linear interpolant [25] connecting the known tendencies to a sample from the initial conditions,
$\mathbf{z}_{\tau}=\tau\mathbf{z}+(1-\tau)\mathbf{z}_{0}.$ (5)
@@
$\mathbf{u}=\mathbf{z}-\mathbf{z}_{0}.$ (6)
@@
However, this velocity is only known during training, and we use a trained neural network as approximation for predictions such that $\mathbf{u}\approx\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z}_{\tau},\mathbf{x}_{t},\mathbf{F}_{t:t+12\,\mathrm{h}},\tau)$} should hold.
@@
Based on variational flow matching [44], the commonly used loss function is reparametrised to account for an appropriate bounding of the sea-ice states with $\mathbf{x}^{L}$ as lower bound and $\mathbf{x}^{U}$ as upper bound. For this reparametrisation, we introduce a new per-variable standard deviation $\mathbf{s}$ which specifies the expected uncertainty of the neural network’s velocity approximation and is used as learnable weighting [61]. As the true flow is a linear interpolant, we can make a mean-field approximation, factorising the loss function over the $n_{\mathrm{vars}}=6$ variables and $n_{\mathrm{grid}}$ unmasked grid points,
@@
$\mathcal{L}_{\mathrm{GenSIM}}(\boldsymbol{\theta},\mathbf{s})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\sum^{n_{\mathrm{vars}}}_{k=1}\sum^{n_{\mathrm{grid}}}_{l=1}J(\boldsymbol{\theta},\mathbf{s},k,l)\right].$ (7)
@@
Using a censored Gaussian as variational distribution [34], we obtain as component cost function,
@@
$J(\boldsymbol{\theta},\mathbf{s},k,l)=$~ $\mathrm{I}(x^{L}_{k}<x_{t+12\,\text{h},k,l}<x^{U}_{k})\left[\frac{1}{2\cdot(s_{k})^{2}}(u_{k,l}-\widehat{v}_{\boldsymbol{\theta},k,l})^{2}+\log(s_{k})\right]$ (8)
@@
$-\mathrm{I}(x_{t+12\,\text{h},k,l}=x^{L}_{k})\left[\log\Phi\Big{(}\frac{u_{k,l}-\widehat{v}_{\boldsymbol{\theta},k,l}}{s_{k}}\Big{)}\right]$

### v2

@@ Abstract
Averaged for 2015–2018, GenSIM’s large-scale patterns in the thermodynamics and dynamics closely resemble those of the coupled neXtSIM-OPA simulation (Fig. 4). Thermodynamically, GenSIM exhibits the same regions of ice growth and melting as neXtSIM (Fig. 4, a and b): major growth occurs in the Laptev sea and Kara sea, as well as the Canadian archipelago and central Arctic, while the strongest melt is found around Greenland, in the Hudson bay, the Barents sea, and the Chukchi sea. The melt-dominated regions are linked to bottom melting (Fig. 4, c), as sea ice is exposed to warmer water from the Atlantic or Pacific ocean therein.
These spatial patterns by GenSIM cannot be explained by the average 2-metre temperature alone (Fig. 4, d). Instead, GenSIM infers them from combining all atmospheric forcings together with the land mask, which allows it to reproduce signatures that in nature arise from ocean–ice interactions. As a result, the neural network identifies the same key melting zones and yields melting rates comparable to those of neXtSIM-OPA. GenSIM implicitly captures the long-term thermodynamical impact of ocean-ice interactions.
@@
Although the sea-ice dynamics are dominated by the atmosphere on daily scales (?), the dynamics of sea ice and ocean remain aligned on year-long scales, as they are both driven by the general circulation from the atmosphere. Consequently, we find an agreement in the sea-ice drift and the ocean surface velocity in coupled sea-ice–ocean simulations (Fig. 4, f and g): sea ice circulates in front of the Alaskan and Canadian coast due to the Beaufort gyre, while the transpolar drift transports sea ice from growing zones to melting zones through the Fram strait. GenSIM’s sea-ice drift matches these major ocean currents in the Arctic basin (Fig. 4, e), even though the effective Beaufort gyre is displaced. As only forcing available for sea-ice motion, the general atmospheric circulation (Fig. 4, h) determines GenSIM’s overall smoother patterns. Remaining differences can be attributed to the mechanical response of sea ice on shorter time scales and to other indirect factors, such as the sea-ice concentration. Once again, GenSIM has apparently learned the long-term link between coupled sea-ice–ocean dynamics and atmospheric forcings by extracting spatial patterns from these forcings.
@@
Without direct access to external oceanic forcings, GenSIM has nonetheless learned the long-term interplay between sea ice and ocean present in the coupled neXtSIM-OPA simulation. The learning of this interplay emerges from patterns that are encoded in the atmospheric forcings and their relationship to the land mask. Consequently, although trained solely trained to predict for a 12-hour lead time, GenSIM correctly infers slow multi-year effects and uncovers hidden influences from the ocean, even when their signal-to-noise ratio is diminishing for the training task. These results explain how GenSIM can match its long-term evolution to advanced physics-based models and satellite observations in a warming climate.
Consequences for AI-based Earth system models
@@
Our introduced generative sea-ice model, GenSIM, improves upon deterministic models by realistically capturing the full image of the Arctic sea ice: all key indicators are predicted with an unprecedented physical consistency across spatial scales from 10 km to the Arctic basin and temporal scales from 12 hours to multi-decadal climate trends. These findings pave a way towards the use of generative diffusion and flow matching models for surrogates of other Earth system components than the atmosphere. Furthermore, large ensembles produced by such generative models can especially be useful for sea-ice data assimilation (?), which has been hindered by high computational costs and the unavailability of adjoints (?). Hence, generative surrogate models unlock a large potential to improve sea-ice forecasting in the future.
@@
We built upon the observation that generative flow models work similarly to physics-based numerical models: they integrate a dynamical system that is defined by the neural network. This similarity allows us to bridge methods used for physics-based models to generative surrogates. Here, we use domain decomposition to efficiently apply such surrogates at global scale. Combined with the employed transformer architecture, the surrogate is flexible in its use with applications from regional to pan-Arctic scales, generalising across scales (Supplementary Text, figure S9). Thanks to this flexibility, we can train generative surrogates across heterogenous datasets and resolutions (?), potentially further boosting their performance.
@@
GenSIM realistically simulates decade-long climate conditions, suggesting that generative surrogates can generalise sub-daily training objectives to long-term projections. We partially attribute its stability to GenSIM’s physical consistency. The underlying physics manifests in the system’s states as evolving correlations across space, time, and between variables. Consequently, this physical consistency is a strength of generative diffusion and flow matching models, which excel in learning high-dimensional, multi-scale correlations. Moreover, on longer time scales, sea ice is dominantly driven by the thermodynamics and ice-ocean interactions. With both well-captured when atmospheric forcings are provided, this further constrains and stabilises the model. Therefore, the long-term stability is likely an inherently emerging capability of generative flows employed as sea-ice models.
@@
The successful extrapolation to long-term evolutions from short-term tasks emerges from learning the influence of slow components on the faster, targetted, time scales. This allows us for example to unveil the interactions of the ice with the ocean, although GenSIM is solely forced by the atmosphere. Consequently, generative flow models can retrieve hidden dynamics in the data, even for weak signal-to-noise ratios.
@@
The most common strategy for Earth system modelling is to have models for each Earth system component that interact through coupling. Yet, this strategy can fail for AI-based component models: as we have demonstrated, the dynamics of these models show an implicit entanglement between components, here sea ice and ocean. This makes it challenging to isolate the dynamics of a single component due to external forcings from other components. Therefore, our findings suggest the need to rethink how to do coupled Earth system modelling, when using AI-based models.
Figure 1: Schematic of the efficient generative AI-based sea-ice model GenSIM. Based on flow matching, GenSIM jointly predicts all important sea-ice states for a lead time of 12 hours. (a) Conditioned on initial states $\mathbf{x}_{t}$ and external atmospheric forcings $\mathbf{F}_{t:t+12\,\text{h}}$, the model iteratively generates a prediction $\widehat{\mathbf{x}}_{t+12\,\text{h}}$ 12 hours later. (b) Within an iterative flow step, at arbitrary pseudo time $s\in[0,1]$, the full Arctic domain is decomposed into overlapping subdomains and processed in parallel with a scale-aware transformer, which predicts a velocity, used to update the fields to the next pseudo tome step $s^{\prime}\geq s$.
Figure 2: Physical consistency of GenSIM. Snapshots from a 30-year-long GenSIM simulation for (a–c) 2024-02-13 03:$0$ UTC and (d–f) 2024-02-13 15:$0$ UTC with shown sea-ice concentration (a&d), divergence rate (b&e), and change in the sea-ice thickness within 12 hours (c&f). Regions of interest for the (I) opening of leads, (II) healing, and (III) ridging are shown in black boxes.
Figure 3: Overview over GenSIM’s long-term projection qualities. (a) The simulated monthly averaged sea-ice volume in April (maximum) and September (minimum) for GenSIM, neXtSIM-OPA, PIOMAS, the merged CS2/SMOS product, and IceSAT2. Note that CS2/SMOS and IceSAT2 are unavailable for September. (b) The seasonal cycle of the average volume averaged for the period 2015–2018 with the same colouring as for (a). (c, d) The change in the winter thickness between 2020-2024 to 2000–2004 for GenSIM and PIOMAS. (e, f) The winter sea-ice thickness averaged between 2015–2018 for GenSIM and neXtSIM. We define the winter season as January to April.
Figure 4: GenSIM exhibits long-term ocean-like patterns. (a–d) Volume change within 12 hours for (a) GenSIM, (b) neXtSIM, (c) net bottom growth - bottom melt from neXtSIM, and (d) average 2-metre temperature from the ERA5 reanalysis, averaged for 2015–2018. The black contour line indicate the $1.8\,^{\circ}\text{C}$ isotherm from the averaged 2-metre temperature. The dynamical component due to advection has been removed from the changes in (a) and (b). (e–h) Speed (background) and streamlines (black lines) for (e) sea ice in GenSIM, (f) sea ice in neXtSIM, (g) ocean surface from the neXtSIM-OPA simulation, and (h) the 10-metre wind from the ERA5 reanalysis, averaged for 2015–2018.
Acknowledgments
We would like to thank Lorenzo Zampieri, Joffrey Dumont Le Brazidec, and Steffen Tietsche (ECMWF), through discussing they helped to focus this study. Additionally, we acknowledge Guillaume Boutin (NERSC) for providing the training dataset and other members from the SASIP project which gave helpful comments along the way. CEREA is a member of the Institut Pierre-Simon Laplace (IPSL).
Funding:
This study is a contribution to the SASIP project funded under Grant no. G-24-66154 by Schmidt Science–a philanthropic initiative that seeks to improve societal outcomes through the development of emerging science and technologies. TSF, MB, CD, and AF additionally received financial support from INSU/CNRS by the project DeepGeneSIS (PNTS). This work was granted access to the HPC resources of IDRIS under the allocations 2021-AD011013069, 2022-AD011013069R1, 2023-AD011013069R2, and 2024-AD011013069R3 made by GENCI.
Author contributions:
Pierre Rampal (PR), Alberto Carrassi (AC), and Marc Bocquet (MB) acquired the project. Tobias Sebastian Finn (TSF), MB, and Alban Farchi (AF) initialised the scientific questions based on previous work of Charlotte Durand (CD) and Flavia Porro (FP). TSF, MB, and PR refined these scientific questions and prepared an analysis strategy. TSF and CD organised the data. TSF created the codebase and performed the experiments. TSF, MB, and PR analysed and discussed the results. TSF wrote and revised the paper, with MB, PR, CD, FP, AF, and AC reviewing.
Competing interests:
There are no competing interests to declare.
Data and materials availability:
The code for GenSIM and its training is available at https://github.com/cerea-daml/gensim, while the trained model checkpoints can be found at https://huggingface.co/tobifinn/gensim. Scripts to produce the training data are included in the code. The sea-ice model data used for training based on work within the SASIP project (?). The atmospheric forcing data is based on the ERA5 reanalysis (?) and available from the Copernicus Climate Change Service (C3S) Climate Data Store (CDS). This work contains data modified from Copernicus Climate Change Service information. Neither the European Commission nor ECMWF is responsible for any use that may be made of the Copernicus information or data it contains.
Supplementary materials
Materials and Methods
Supplementary Text
Figures S1 to S11
References (52-0)
Supplementary Materials for
Generative AI models capture realistic sea-ice evolution from days to decades
Tobias Sebastian Finn, Marc Bocquet, Pierre Rampal
Charlotte Durand, Flavia Porro, Alban Farchi, Alberto Carrassi
Corresponding author. Email: tobias.finn@enpc.fr
This PDF file includes:
Materials and Methods
Supplementary Text
Figures S1 to S11
Materials and Methods
Data used for GenSIM
@@ Acknowledgments
The model is trained and compared to a 24-year-long simulation (1995-2018) (?) produced by neXtSIM (?), an advanced sea-ice model employing Lagrangian finite elements and a brittle Bingham-Maxwell rheology (?) (data available at https://github.com/sasip-climate/catalog-shared-data-SASIP/blob/main/outputs/OPA-neXtSIM_CREG025-ILBOXE140.md, last access: 18th August 2025). For this simulation, neXtSIM has been coupled to OPA, the ocean component of the NEMO modelling framework (?) and has been forced with the ERA5 reanalysis (?) for the atmosphere. Although neXtSIM-OPA is initialised from a climatology, we use all the data from the start as our goal is to learn a surrogate that emulates the original system as close as possible. The curvilinear mesh of our model corresponds to the Eulerian Arakawa-C mesh of OPA, to which the model output of neXtSIM has been interpolated. GenSIM predicts sea-ice thickness, concentration, damage, $u$- and $v$-drift, and the snow-on-ice thickness. All variables but the damage are averaged in a six hour window, while the damage is an instantaneous state.
@@
In addition to the initial conditions, GenSIM takes external atmospheric forcings from the ERA5 reanalysis dataset (?). Originally on a latitude-longitude mesh with a $1/4^{\circ}$ resolution, the dataset is interpolated to our curvilinear mesh by nearest neighbour. The forcings consist of the 2-metre temperature, the 2-metre specific humidity, and the 10-metre equatorial and meridional wind. The specific humidity is estimated as product of the 2-metre temperature, the 2-metre dew point, and the surface pressure.
@@
To obtain neural network inputs that contain a memory about the past development without having access to the ocean, we construct temperature degree day features from the 2-metre temperature (?, ?). Using a temperature threshold of $T_{\textrm{thres}}=271.35\,\text{K}$ (freezing point at an ocean salinity of about $32\,\text{PSU}$), we construct the positive, $\mathrm{PDD}$, and negative degree days, $\mathrm{NDD}$, by
$\mathrm{PDD}(T_{\mathrm{2m}},\Delta t)$ $=\int^{t}_{t-\Delta t}\mathrm{maximum(T_{\mathrm{2m}}(s)-T_{\textrm{thres}},0)}{\rm d}s$
$\text{and }\mathrm{NDD}(T_{\mathrm{2m}},\Delta t)$ $=\int^{t}_{t-\Delta t}\mathrm{minimum(T_{\mathrm{2m}}(s)-T_{\textrm{thres}},0)}{\rm d}s,$
@@
We evaluate our long-term simulation with the sea-ice–ocean reanalysis product from the Pan-Arctic Ice-Ocean Modeling and Assimilation System (PIOMAS) (?) (data available at https://psc.apl.uw.edu/research/projects/arctic-sea-ice-volume-anomaly, last access: 18th August 2025). The dataset used here is on a one-degree latitudinal-longitudinal mesh and consists of monthly averaged sea-ice thickness and concentration values.
@@
The sea-ice volume is additionally evaluated with the CS2/SMOS product (?) (Version 2.6, data available at https://spaces.awi.de/spaces/CS2SMOS/overview, last access: 18th August 2025), which merges satellite observations from the CryoSat2 and Soil Moisture and Ocean Salinity (SMOS) missions. As CS2/SMOS infers solely the sea-ice thickness, it is complemented by sea-ice concentration observations of the EUMETSAT’s climate data record from the OSISAF project (?), which merges satellite measurements from SSMR, SSM/I, and SSMIS. Both datasets are on the EASE mesh with $25\,\text{km}$ resolution (Identifier: OSI-450-a1 and OSI-430-a, Version 3, data available at https://osi-saf.eumetsat.int/products/sea-ice-products, last access: 5th November 2025).
Also, we use the thickness satellite product of IceSAT2 (?, ?) (Version 3, data available at https://nsidc.org/data/is2sitmogr4/versions/3, last access: 5th November 2025) for sea-ice volume validation. This thickness data is complemented by the sea-ice concentration from the NOAA/NSIDC climate data record (?) (Version 5, data available at https://nsidc.org/data/g02202/versions/5, last access: 5th November 2025).
Surrogate modelling with censored flow matching
@@
In the following, we expose the most important equations to understand how GenSIM is trained. For a more rigorous treatment and a mathematical derivation of the equations, we refer to Supplementary Text Mathematical framework for surrogate modelling.
@@
For a forecast window of 12 hours, based on the current state $\mathbf{x}_{t}$ and the forcings in the same window $\mathbf{F}_{t:t+12\,\mathrm{h}}$, the surrogate model predicts the sea-ice state $\mathbf{x}_{t+12\,\mathrm{h}}$–consisting of sea-ice thickness, concentration, damage, the drift in $x$- and $y$-direction, and the snow-on-ice thickness. Instead of predicting the future state directly, the model outputs predicted tendencies $\widehat{\mathbf{z}}_{1}$, which are scaled by the per-variable climatological standard deviation of the tendencies, $\boldsymbol{\sigma}_{\textrm{clim}}$, such that we can recover the prediction by
@@
$\widehat{\mathbf{x}}_{t+12\,\text{h}}=\mathbf{x}_{t}+\boldsymbol{\sigma}_{\textrm{clim}}\odot\widehat{\mathbf{z}}_{1}.$ (S1)
@@
In our autoregressive framework, the subsequent prediction $\widehat{\mathbf{x}}_{t+24\,\text{h}}$ is performed with the previous prediction and updated forcings $\mathbf{F}_{t+12\,\mathrm{h}:t+24\,\mathrm{h}}$ as new inputs. This way, we can extend the prediction as long as possible and desired (?, ?).
@@
For training, we have access to a dataset with data triplets $(\mathbf{x}_{t},\mathbf{x}_{t+12\,\mathrm{h}},\mathbf{F}_{t:t+12\,\mathrm{h}})\sim\mathcal{D}$. From these data triplets, we can construct the targetted tendencies, $\mathbf{z}_{1}=(\mathbf{x}_{t+12\,\mathrm{h}}-\mathbf{x}_{t})/\boldsymbol{\sigma}_{\textrm{clim}}$.
@@
For GenSIM, we instantiate a generative flow to produce stochastic predictions (?, ?). The generative flow is defined by the integration of an ordinary differential equation (ODE) in pseudo time $s\in[0,1]$,
$\frac{\textrm{d}\mathbf{z}_{s}}{\textrm{d}s}=\mathbf{u}$ (S2)
@@
with its generative velocity $\mathbf{u}$ and the initial noise sample at pseudo time $s=0$,
$\mathbf{z}_{0}\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}).$ (S3)
@@
Note that the subscript $s$ describes the dependency on the pseudo time, whereas $t$ is the real time.
@@
During training, when we know the targetted tendencies, we construct a linear interpolant (?) connecting the known tendencies to an initial noise sample,
$\mathbf{z}_{s}=s\mathbf{z}_{1}+(1-s)\mathbf{z}_{0}.$ (S4)
@@
$\mathbf{u}=\mathbf{z}-\mathbf{z}_{0}.$ (S5)
@@
However, this velocity is only known during training, and we use a trained neural network as approximation for predictions such that $\mathbf{u}\approx\widehat{\mathbf{v}}_{\boldsymbol{\theta}}(\mathbf{z}_{s},\mathbf{x}_{t},\mathbf{F}_{t:t+12\,\mathrm{h}},s)$} should hold.
@@
Based on variational flow matching (?), the commonly used loss function is reparametrised to account for an appropriate bounding of the sea-ice states with $\mathbf{x}^{L}$ as lower bound and $\mathbf{x}^{U}$ as upper bound. For this reparametrisation, we introduce a new per-variable standard deviation $\boldsymbol{\sigma}$ which specifies the expected uncertainty of the velocity approximation and is used as learnable weighting (?). As the true flow is a linear interpolant, we can make a mean-field approximation, factorising the loss function over the $n_{\mathrm{vars}}=6$ variables and $n_{\mathrm{grid}}$ unmasked grid points,
@@
$\mathcal{L}_{\mathrm{GenSIM}}(\boldsymbol{\theta},\mathbf{s})=\mathbb{E}_{\tau,\mathbf{z}_{0},\mathcal{D}}\left[\sum^{n_{\mathrm{vars}}}_{k=1}\sum^{n_{\mathrm{grid}}}_{l=1}J(\boldsymbol{\theta},\boldsymbol{\sigma},k,l)\right].$ (S6)
@@
Using a censored Gaussian as variational distribution (?), we obtain as component cost function,
@@
$J(\boldsymbol{\theta},\mathbf{s},k,l)=$~ $\mathrm{I}(x^{L}_{k}<x_{t+12\,\text{h},k,l}<x^{U}_{k})\left[\frac{1}{2\cdot(\sigma_{k})^{2}}(u_{k,l}-\widehat{v}_{\boldsymbol{\theta},k,l})^{2}+\log(\sigma_{k})\right]$ (S7)
@@
$-\mathrm{I}(x_{t+12\,\text{h},k,l}=x^{L}_{k})\left[\log\Phi\Big(\frac{u_{k,l}-\widehat{v}_{\boldsymbol{\theta},k,l}}{\sigma_{k}}\Big)\right]$